# Diffusion in Zero-Shot Learning for Environmental Audio

### v1

@@ II-C Environmental Audio Zero-Shot Learning
Limited research exists for environmental audio zero-shot learning. One of the earliest works proposes a Twin neural network [66], tested on a small dataset containing ten classes. Xie and Viranen [18] used a more substantial dataset of fifty classes. In addition to a bilinear compatibility model [67] that created a baseline for the ESC-50 dataset, they proposed two partition strategies for five-fold cross-validation. These partitions are used in subsequent works in the field to allow researchers to compare methods reliably.
Further work in the field [68, 69] brought new state-of-the-art results and critical considerations when training zero-shot networks. Xie and Viranen [68] highlight the importance of the unseen and seen class distinction, where they identify that VGGish pre-trained on AudioSet is unsuitable for zero-shot learning on ESC-50 due to the presence of unseen classes in AudioSet.
AudioSet contains embeddings rather than audio clips. Embeddings are a critical aspect of zero-shot learning methods, and some research works explore improving these embeddings [19]. Furthermore, it is vital to maintain the distinction between ‘seen’ and ‘unseen’ classes. Therefore, training from raw data is essential. AudioSet provides the location of the YouTube videos the clips come from, but it is against the Terms of Service to download the raw data. AudioSet is, therefore, unsuitable for zero-shot learning research.
@@
Some research focuses on improvements through embeddings. While not often the main focus of methods, they are the foundations of the zero-shot method. The remaining pipeline is limited to the information that the embeddings can capture. The methods in this paper use work by Sims et al. [19] to improve the performance through embeddings by adding synonyms.
III Method
@@ III-A Zero-Shot Learning
The zero-shot learning problem is defined as the following. There exists a set $X$ of seen data and a set $Y$ of unseen data such that $X\cap Y=\varnothing$. The zero-shot learning method aims to learn a function
$f(z,C)=c$ (1)
where $z\in X\cup Y$ is a data sample; $C$ is the set of all classes; and $c\in C$ is the class that the sample $z$ comes from. After training a model to learn $f$ using all samples in $X$, the model is tested on all samples in $Y$.
The literature frequently divides zero-shot learning methods into three components. One component is the auxiliary data function $\sigma(c)$ that takes a class $c\in C$. The literature overwhelmingly uses word embedding models or human-engineered attributes for $\sigma$. $\sigma$ changes a class label into high-level information.
A second component is the feature embedding function $\tau(z)$, where $z\in X\cup Y$ is any feature data in the dataset. The feature data in the case of environment audio is audio clips. $\tau$ will convert the feature data into high-level information, similarly to $\sigma$. The leading method for $\tau$ in the literature involves training a traditional supervised classification model on the seen classes $Y$ and taking the second last layer as an information-rich embedding layer. In computer vision, a popular method is to use ResNet, pre-trained on ImageNet.
The third component ties the previous two together to retrieve a class prediction. This component learns a function
$\xi(\tau(z),\sigma(C))=c$ (2)
that takes the high-level information output of the other two components and outputs the corresponding class $c$ for a sample $z$. This component of zero-shot learning is explored heavily across the literature.
@@ III-B Embeddings
An embedding is an encoded form of data that is easy for a machine learning model to process. They are typically information-rich low-dimensional vectors representing data that would be cumbersome and computationally intensive in their raw form.
@@
The $\sigma$ function is the Word2Vec [60] pre-trained model. This model takes class labels and outputs an embedding vector for that word. The space that Word2Vec has learned is information-rich, with semantically similar words existing closer together in the space than semantically distant words. Word2Vec gives an embedding in $\mathbb{R}^{300}$.
We use work by Sims et al. [19] to enrich the word embeddings with synonyms. The average word embedding of the class label and synonyms are used as auxiliary information. For ESC-50, the synonyms from the original paper were used. FSC22 has not been used in previous literature on synonyms, and so we human-engineer the synonyms, where overlapping classes use the same synonym list from ESC-50.
To learn $\tau$ and obtain audio embeddings, we train a modified YAMNet model, shown to perform better than other methods [19], on the seen class data. As shown in Figure 2, the second-last layer of the model is a fully connected embedding layer of size 128, designed to capture high-level semantic information for use as an audio embedding vector.
Thus both $\sigma$ and $\tau$ have been defined, where the output of $\sigma(c)$ is a class embedding of dimension $300$ for the class $c$, and $\tau(w)$ is an audio embedding of dimension $128$ for the audio data $w\in X\cup Y$.
@@ III-C ZeroDiffusion
Figure 2: ZeroDiffusion involves training three models. The first is the audio embedding model, used to retrieve audio embeddings. Seen classes are used to train this model. Next, the diffusion model is trained using audio embeddings from the first model and word embeddings. This process is done using seen classes. Finally, a compatibility model is trained using seen data and synthetic unseen data generated using the trained diffusion model.
Diffusion involves training a model to reconstruct data from noise. During training, the input to the model is a data sample with noise added, and the expected output is the original data sample. The amount of noise added to the data increases as training progresses, where the first few epochs have very little noise added to the data and the final epochs result in mostly or entirely noise. This process gradually trains a model to generate data from noise.
@@
A diffusion model conditioned on auxiliary data is trained with the input to the model being paired embeddings $\sigma(c)$ and $\tau(w)$ where $c$ is the corresponding class for the audio sample $w\in X$. The model learns to recreate the audio embedding $\tau(w)$ based on the given class embedding $\sigma(c)$.
@@
The audio embedding $\tau(w)$ is mixed with noise from a normal distribution with mean 0 and variance 0.1 to create a gradually noisy embedding vector
$\tilde{\tau}(w)(p)=\tau(w)+p\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,0.1),$ (3)
where $p$ is the epoch percentage.
@@
The input to the network is the noise-injected audio embedding $\tilde{\tau}(w)$ concatenated with the class embedding $\sigma(c)$. The first embedding has a dimension of 128, and the second has a dimension of 300, resulting in an input layer of size 428. Due to the concise and information-rich nature of embeddings, the diffusion architecture is small and uses fully connected layers. The first hidden layer is a fully connected layer of size 128, to reduce the dimension to the expected output size. This is followed by batch normalisation for training stabilisation and then a Leaky ReLU activation to add a non-linearity component. Following Leaky ReLU, there is a fully connected output layer of size 128 and finally a hyperbolic tangent activation, which bounds the output to the expected range for the audio embedding space. The network output is an embedding of size 128 to match the size of the expected recreated audio embedding.
The loss is the mean squared error between the embedding the model is trying to recreate and the output of the model. During training, gradients are clipped to a maximum norm of 1 to stabilise training. An Adam optimiser is used.
@@
Our new diffusion model is trained on seen class data. The aim is for the model to generalise to generating unseen data. Unlimited synthetic data for the unseen classes can be generated using the unseen class embeddings and random noise. After creating an unseen dataset in this manner, a bilinear compatibility network [67, 18] is trained with both the true seen data and synthetic unseen data. The bilinear network uses a cross-entropy loss function and an Adam optimiser.
@@ IV Experiments
Experiments are performed on the ESC-50 and FSC22 datasets. The state-of-the-art in zero-shot learning environmental audio, the compatibility bilinear method [67, 18] is compared with existing methods from computer vision [14, 15], and our new method ZeroDiffusion.
ESC-50 is split into five partitions of ten classes. One partition is held as the test set. The remaining four partitions are used for four-fold cross-validation, where three of the partitions are used in training and the remaining one is used for validation. The validation set is used to tune the model and its hyperparameters before using the test set. This design is modified from the five-fold cross validation used in previous work [68, 19] to include a held-out test set. Using a four-fold strategy for hyperparameter tuning allows for more generalised hyperparameter tuning and a better understanding of the performance of the model. The final experiments on the held-out test set give a good indication of how the model performs on a completely unseen set of classes and data. ESC-50 is a suitable dataset for this strategy as it is composed of many classes, allowing for each partition to have ten classes. The dataset is correctly labelled and clean, making it ideal for testing.
FSC22 is split into three sets, a test set with 7 classes, a validation set of 7 classes, and a training set with the remaining 13 classes. The validation set is used for hyperparameter tuning, and then the method is evaluated using the test set. This dataset has almost half the number of classes compared to ESC-50, reducing the amount of flexibility in forming a strategy. Additionally, the number of classes is uneven and has few divisors. If it were partitioned into nine lots of three, the number of classes in each set would be extremely small for zero-shot learning while also creating a lot of folds to manage. If three lots of nine were used, the training set would be very small compared to the usual split in zero-shot learning, where there are more training classes than test classes. Therefore, the validation and test sets are chosen to be the same size for consistency, and train with a larger number, being almost the size of the test and validation sets combined. Therefore, $k$-fold validation is not performed for this dataset. The dataset does overlap with the classes in ESC-50, but there are new classes not present in ESC-50. Some classes have questionable labels, including the class ‘silence’, which appears to contain a variety of soundscapes including birds chirping and rain falling, which are classes elsewhere in the dataset. Because the dataset is not entirely well-labelled, this is considered a harder dataset compared to ESC-50.
@@
All experiments are repeated ten times with the average and standard deviation of the accuracy recorded. The experiments were run on an NVIDIA GeForce RTX 3080 Ti Laptop GPU with 16GB of video memory.
Index Class Index Class Index Class
0 Dog 17 Pouring water 34 Can opening
1 Rooster 18 Toilet flush 35 Washing machine
2 Pig 19 Thunderstorm 36 Vacuum cleaner
3 Cow 20 Crying baby 37 Clock alarm
4 Frog 21 Sneezing 38 Clock tick
5 Cat 22 Clapping 39 Glass breaking
6 Hen 23 Breathing 40 Helicopter
7 Insects (flying) 24 Coughing 41 Chainsaw
8 Sheep 25 Footsteps 42 Siren
9 Crow 26 Laughing 43 Car horn
10 Rain 27 Brushing teeth 44 Engine
11 Sea waves 28 Snoring 45 Train
12 Crackling fire 29 Drinking, sipping 46 Church bells
13 Crickets 30 Door knock 47 Airplane
14 Chirping birds 31 Mouse click 48 Fireworks
15 Water drops 32 Keyboard typing 49 Hand saw
16 Wind 33 Door, wood creaks
TABLE I: Class indices and names in ESC-50. This follows the convention from the dataset.
Index Class Index Class Index Class
0 Fire 9 Axe 18 Footsteps
1 Rain 10 Chainsaw 19 Clapping
2 Thunderstorm 11 Generator 20 Insect
3 Water Drops 12 Handsaw 21 Frog
4 Wind 13 Firework 22 Bird Chirping
5 Silence 14 Gunshot 23 Wing Flapping
6 Tree Falling 15 Wood Chop 24 Lion
7 Helicopter 16 Whistling 25 Wolf Howl
8 Vehicle Engine 17 Speaking 26 Squirrel
TABLE II: Class indices and names in FSC22. While the same order is followed from the dataset, here the classes are zero-indexed rather than one-indexed.

### v2

@@ II-C Environmental Audio Zero-Shot Learning
Limited research exists for environmental audio zero-shot learning. An early study proposes a Twin neural network [46], tested on a small dataset containing ten classes. A Twin neural network is an architecture with two identical networks that process two inputs at the same time to learn a similarity function. Xie and Virtanen [19] used a more substantial dataset of fifty classes. They introduce a bilinear compatibility network, which builds on from Attribute Label Embedding (ALE) [31]. A compatibility function is learnt between the audio data and class embeddings by computing:
$f(x,y)=x^{T}Wy$ (1)
Where $x$ is the audio embedding, $y$ is the class embedding, and $W$ is a trainable weight matrix. This maps the inputs to a shared latent space for comparison directly, without the use of generative methods.
This method sets an initial baseline for the ESC-50 dataset, and the authors introduce two partition strategies for five-fold cross-validation. These partitions are used in subsequent works in the field to allow researchers to compare methods reliably.
Further work in the field [47, 48] brought new state-of-the-art results and critical considerations when training zero-shot networks. Xie et al. [47] highlight the importance of the unseen and seen class distinction when using pre-trained models. A VGGish model trained on AudioSet (a large environmental audio dataset) may produce rich embeddings, but a majority of classes in ESC-50 occur in AudioSet. The classes in ESC-50 which are meant to be unseen would have instead been seen by the VGGish network during pretraining on AudioSet. This breaks the zero-shot learning formulation. It is important to understand that it is the feature data that should be unseen - the classes are known, and therefore a pre-trained word embedding model does not invalidate the method as zero-shot learning.
Wu et al. use contrastive audio-text representation learning with CLAP to perform zero-shot learning on multiple classes[49]. However, like other CLAP and CLIP-based methods, these models require training data from large datasets in the target modality and this is not always feasible in a given domain. Additionally, even though repeated samples across datasets are removed, whole classes aren’t removed. This is evident in the ESC-50 and FSD50k case, where the text reports 630 overlapping files, even though within one overlapping class, ‘dog’, there are over 900 samples in FSD50K that would need to be removed. Indeed, this would result in extremely few classes remaining since audio sources for datasets are very limited and many datasets are from YouTube or Freesound, with many of the same classes appearing across audio datasets. However, this inflates the reported zero-shot classification performance.
@@
Some research focuses on improvements through embeddings. While not often the main focus of methods, they provide the foundations for unseen classification. Any embedding-based method is limited to the information that the embeddings can capture. The methods in this paper use work by Sims et al. [20] to improve the performance through embeddings by adding synonyms, except ARCA23K-FSD (due to size) and GTZAN (lack of synonyms for genres).
II-D Contributions and Novelty
This section will detail the contributions of the presented paper, in the context of the literature. Literature for zero-shot environmental sound classification is lacking, and no existing literature in the domain explores generative methods. The experiments in this work not only benchmark environmental audio datasets on the existing generative methods CADA-VAE and LisGAN but also introduce a novel diffusion-based method that is not tied to a particular modality. Code for the novel ZeroDiffusion method is provided at https://github.com/ysims/ZeroDiffusion.
Two publications exist in zero-shot learning using diffusion ([44, 45]); however, they both use Stable Diffusion directly to classify unseen data, with no further training of any model. Their methods are specific to computer vision and textual class labels through their use of Stable Diffusion. Meanwhile, the novel diffusion method presented in this work is modality-agnostic, because it uses embeddings as input to the diffusion model. Any class embedding can be used to condition the diffusion model, and any feature embedding can be used for training, regardless of what modality it comes from. The method involves training a diffusion model, which is used to generate unseen embeddings to train a compatibility-based classifier. This generative-style zero-shot learning method is significantly different to the previous work on Stable Diffusion and provides greater flexibility.
In addition to designing the first modality-agnostic diffusion method for zero-shot learning, this work additionally performs the first zero-shot benchmarks of FSC22, ARCA23K-FSD and TAU Urban Acoustic Scenes 2019.
III Methods
@@ III-A Zero-Shot Learning
The zero-shot learning method aims to learn a similarity function $f(w,z)$, where $w$ is a feature data sample and $z\in X\cup Y$ is a class. $X$ is the set of unseen classes and $Y$ is the set of seen classes. $X\cup Y$ is the set of all classes and $X\cap Y=\varnothing$. After training a model to learn $f(w,z)$ using all samples belonging to classes in $X$, the model is tested on all samples belonging to classes in $Y$.
Many zero-shot learning methods can be considered to contain three distinct components. One component is the auxiliary data function $\sigma(z)$ that takes a class $z\in X\cup Y$. The literature overwhelmingly uses word embedding models or human-engineered attributes for $\sigma$. $\sigma$ changes a class label into high-level information.
A second component is the feature embedding function $\tau(w)$, where $w$ is any feature data in the dataset. The feature data in the case of environment audio is audio clips. $\tau$ will convert the feature data into high-level information, similarly to $\sigma$. The leading method for $\tau$ in the literature involves training a traditional supervised classification model on the seen classes $Y$ and taking the second last layer as an information-rich embedding layer.
The third component ties the previous two together to retrieve a class prediction. This component learns a similarity function $\xi(\tau(w),\sigma(z))$ that takes the high-level output of the two embedding components and outputs the similarity of $w$ and $z$. By comparing $w$ against all classes ($X$ during training and $Y$ during testing), either top-1 or thresholding can be used to retrieve the prediction.
@@ III-B Embeddings
An embedding is an encoded form of data that is simple for a machine learning model to process. They are typically information-rich low-dimensional vectors representing data that would be cumbersome and computationally intensive in their raw form.
@@
The $\sigma(z)$ function in this work is the Word2Vec [33] pre-trained model. This model takes class labels and outputs an embedding vector for that word. The space that Word2Vec has learned is information-rich, with semantically similar words existing closer together in the space than semantically distant words. Word2Vec gives an embedding in $\mathbb{R}^{300}$.
We use work by Sims et al. [20] to enrich the word embeddings with synonyms. The average word embedding of the class label and synonyms are used as auxiliary information. For ESC-50, the synonyms from the original paper were used. Synonyms are human-engineered for the other datasets, where overlapping classes across datasets use the same synonym lists. ACRA23K-FSD and GTZAN are used without added synonyms, due to size and difficulty in getting synonyms for music genres. Synonyms used are detailed in the supplementary materials.
To learn $\tau(w)$ and obtain audio embeddings, we train a modified YAMNet model [20] on the seen class data, due to the higher accuracy compared to VGGish when used for zero-shot learning. The input to the modified YAMNet model is a log Mel spectrogram with 64 bins. As shown in Figure 2, the second-last layer of the model is a fully connected embedding layer of size 128, designed to capture high-level semantic information for use as an audio embedding vector.
Both $\sigma(z)$ and $\tau(w)$ have been defined, where the output of $\sigma(z)$ is a class embedding of dimension $300$ for the class $z\in X\cup Y$, and $\tau(w)$ is an audio embedding of dimension $128$ for the audio data $w$.
@@ III-C ZeroDiffusion
Figure 2: ZeroDiffusion involves training three models. The first (1) is the audio embedding model, used to retrieve audio embeddings. The model is trained with seen classes. Next, (2) the diffusion model is trained using audio embeddings and word embeddings from seen classes. Finally, (3) a compatibility model is trained using seen data and synthetic unseen data generated using the trained diffusion model.
We introduce our new method ZeroDiffusion, which applies a novel embedding-space diffusion model to generate synthetic unseen data to train a compatibility classifier. While other diffusion methods use large diffusion models to generate feature embeddings, ZeroDiffusion uses feature embeddings as input to learn to generate unseen data. The diffusion model itself is very lightweight because it works within the embedding space.
Diffusion involves training a model to reconstruct data from noise. During training, the model input is a data sample with added noise, and the expected output is the original data sample. The amount of noise added to the data increases as training progresses, where the first few epochs have very little noise added to the data and the final epochs result in mostly or entirely noise. This process gradually trains a model to generate data from noise.
@@
A diffusion model conditioned on auxiliary data is trained where the input to the model is the paired embeddings $\sigma(z)$ and $\tau(w)$ where $z$ is the corresponding class for the audio sample $w$. The model learns to recreate the audio embedding $\tau(w)$ based on the given class embedding $\sigma(z)$.
@@
The audio embedding $\tau(w)$ is mixed with noise from a normal distribution with a mean of 0 and a variance of 0.1 to create a gradually noisy embedding vector:
$\tilde{\tau}(w,p)=\tau(w)+p\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,0.1),$ (2)
where $p$ is the epoch percentage. Similarly the class embedding $\sigma(z)$ is jittered for diversity using
$\tilde{\sigma}(z)=\sigma(z)+\epsilon,\quad\epsilon\sim\mathcal{N}(0,0.1),$ (3)
@@
The input to the network is the noise-injected audio embedding $\tilde{\tau}(w)$ concatenated with the class embedding $\sigma(z)$. The first embedding has a dimension of 128, and the second has a dimension of 300, resulting in an input layer of size 428. Due to the concise and information-rich nature of embeddings, the diffusion architecture is small and uses fully connected layers. The first hidden layer is a fully connected layer of size 128, to reduce the dimension to the expected output size. This is followed by a Leaky ReLU activation for a non-linearity component and then a dropout of $0.3$ to encourage diversity. Following the dropout layer, there is a fully connected output layer of size 128 and finally a hyperbolic tangent activation, which bounds the output to the expected range for the audio embedding space. The network output is an embedding of size 128 to match the size of the expected recreated audio embedding.
The loss term involves five components: a mean squared error for the reconstruction loss, a maximum mean discrepancy for distribution loss, variance loss to encourage diversity, an average mean squared error of generated and real features in a batch for global centroid alignment, and a cosine similarity loss to encourage directional alignment of embeddings. During training, gradients are clipped to a maximum norm of 1 to stabilise training. An Adam optimiser is used.
@@
Our new diffusion model is trained on seen class data. The aim is for the model to generalise to generating unseen data. Unlimited synthetic data for the unseen classes can be generated using the unseen class embeddings and random noise. After creating an unseen dataset in this manner, a non-linear compatibility network [31, 48] is trained with both the true seen data and synthetic unseen data. The compatibility network uses a cross-entropy loss function and an Adam optimiser.
III-D Non-linear ALE
Xie et al. introduce the non-linear version [48] of the bilinear compatibility network [19], which is based on ALE [31]. This method is used as a baseline due to its good performance amongst the zero-shot environmental sound classification literature and its simplicity. The lack of generators (as opposed to GAN and VAE methods) and its single classification nature means that it has few hyperparameters to tune and training across runs is stable, as shown in the experiments presented in section IV.
The method learns a function:
$f(w,z)=(B\cdot(\tanh(A\cdot w)))^{T}\cdot z$ (4)
where $w$ is the audio embedding, $z$ is the class embedding, $A$ and $B$ are both trainable weight matrices, and $\tanh$ is the hyperbolic tangent activation. $f(w,z)$ represents the similarity between $w$ and $z$, and is calculated for every $z$. This creates an encoded vector where the $i$th position in the vector represents the confidence or similarity for the $i$th class. The model is independent of the number of classes, unlike typical classification models where the last layer is equal to the number of classes.
III-E LisGAN
LisGAN [16] is a GAN method that uses ‘soul samples’ to regulate data generation. It trains a conditional Wasserstein GAN [38] where the generator uses class embeddings and noise to generate feature data and the discriminator not only determines whether a sample is generated or real but also has a supervised classification component to guide the conditional training. Each class has a soul sample - a representative sample for the class - and the generator is trained to generate samples for a given class close to that class’s soul sample. For seen classes, the soul sample is the average of that class’s data. For unseen classes, the class embedding defines the soul sample.
The discriminator’s classification output is used to screen synthetic samples. Synthetic samples with high confidence are used to train a final classifier. Additionally, unseen and unlabelled data that the discriminator’s classifier predicts with high confidence are used in training the final classifier. This use of unlabelled unseen data is transductive zero-shot learning; however, in this work, only the inductive case is considered. Therefore, we remove the training on unlabelled seen data. The core idea of the LisGAN method remains, where soul samples are used to guide GAN training, and the discriminator’s classifier is used to filter data to train a final classifier. Despite removing a component of LisGAN, it outperformed other GAN methods (vanilla conditional Wasserstein GAN, f-CLSWGAN) in pilot studies.
III-F CADA-VAE
CADA-VAE [15] is a popular method in computer vision-based zero-shot learning that competes with new methods. In this work, CADA-VAE is applied to the domain of environmental audio to evaluate its performance as a variational autoencoder-based method for zero-shot environmental sound classification. The core of the method is the two variational autoencoders (VAEs) and the loss function. One VAE is designed to encode and decode class embeddings, while the other is designed to encode and decode feature embeddings (typically image embeddings, but in this work, audio embeddings are used). Not only are the VAEs trained in a typical manner with reconstruction loss and Kullback–Leibler divergence, but they are also trained with cross-alignment and distribution alignment. Cross-alignment measures the ability of the decoders to decode data from the other modality into their modality - that is, take an encoded class embedding and decode it into a feature embedding, and vice versa. Distribution alignment measures the difference between the distribution of the class embedding autoencoder and the feature embedding autoencoder. Both of these loss terms aim to create a shared latent space between both modalities. The latent space is then used to train a classifier, which performs zero-shot learning at test time.
@@ IV Experiments
Experiments are performed on the ESC-50, ARCA23K-FSD, FSC22, UrbanSound8K, TAU Urban Acoustic Scenes 2019 and GTZAN datasets. The state-of-the-art in zero-shot learning environmental audio, the non-linear ALE method [31, 48] is compared with existing methods from computer vision [15, 16], and our new method ZeroDiffusion.
@@
All experiments are repeated ten times with the average and standard deviation of the accuracy recorded. The experiments were conducted on an NVIDIA GeForce RTX 3080 Ti Laptop GPU with 16GB of video memory.
IV-A Datasets
Environmental sound classification severely lacks large datasets and diversity. ESC-50, FSC22, ARCA23K-FSD and UrbanSound8k are sourced from Freesound, a repository of user-contributed tagged audio clips released under the Creative Commons license. Some clips in these datasets may overlap, and classes are typically similar. TAU 2019 and GTZAN are collected from other sources, and investigate the problems of acoustic scene recognition and music genre classification, respectively. These datasets add diversity to the experiments both in audio clip source and task.
IV-A1 ESC-50
ESC-50 [51] is a clean environmental sound classification dataset of 50 classes which can be arranged into five balanced categories with ten classes each. There are 40 audio clips per class, each five seconds long. The dataset is from Freesound and has been carefully curated to ensure a clean dataset. It is a popular dataset due to its reasonable size and correct labels. The large number of classes in the dataset motivates using k-fold cross-validation, as done in previous work [19, 20].
IV-A2 FSC22
FSC22 [52] is a sound classification dataset focusing on forest sounds. It has 27 classes, each with 75 5-second-long audio clips. The clips are sampled from Freesound. The dataset is not as clean as ESC-50, particularly the ‘silence’ class where some samples contain the sound of bird calls, despite the existence of a bird class. Because of this, the dataset is considered harder to learn.
IV-A3 ARCA23K-FSD
ARCA23K-FSD [53] is a single-label dataset for environmental audio classification derived from the multi-label FSD50K environmental audio dataset, which in turn is collected from Freesound. The dataset contains 70 unbalanced classes and over 23,000 clips. FSD50K is manually human-labelled and contains clips varying in length from 0.3 seconds to 30 seconds. Due to the large number of classes, a similar approach to ESC-50 can be undertaken where k-fold cross-validation can be used for evaluation.
IV-A4 UrbanSound8k
UrbanSound8k [54] is an environmental sound classification dataset collected from Freesound. Each clip is no more than 4 seconds long, and a total of 8,732 clips are in the dataset. There are 10 unbalanced classes. The small amount of classes is problematic for zero-shot learning, as even splitting into three sets (train, validation and test) results in very little in each partition. An alternative is to use only train and test partitions.
IV-A5 TAU Urban Acoustic Scenes 2019
TAU 2019 [55] is an acoustic scene dataset where the aim isn’t to classify specific sounds but rather the type of location the sound is from, for example, a park or an airport. Each audio segment is 10 seconds long and belongs to one of 10 acoustic scene classes. Each scene has 1,440 audio clips. The audio clips in the dataset were collected by the authors rather than scraped or curated from existing sources online. This positions it as a useful dataset due to its unique audio clips, whereas the Freesound datasets have overlap due to their shared source. Like UrbanSound8k, there are only 10 classes so splitting the data into two sets (train and test) is more reasonable than adding a validation partition.
IV-A6 GTZAN
GTZAN [56] is a music genre dataset of 10 genres with 100 audio files for each class. Each clip is 30 seconds long. While this is an adjacent domain to environmental sound classification, it provides diversity to the experiments when many of the other datasets come from the same source (Freesound). The number of classes is equal to TAU 2019 and UrbanSound8k, motivating the use of the same partitioning strategy.